Your friend wants to start a company and promises that you will earn eight times your investment in the next five years.

What is the quarterly return on your investment?
Answer #1

Present\: value \times \left ( 1+i \right )^{n}=Future\, value

1 \times \left ( 1+i \right )^{4\times 5}=8

\left ( 1+i \right )=8^{\frac{1}{20}}

\left ( 1+i \right )=1.10957

i \right )=0.10957

Quarterly return is 10.957%.